A bullet is fired from a gun. The force on the bullet is given by


F = 600 - 2 × 105 t


Where, F is in newtons and t in seconds. The force on the bullet becomes zero as soon as it leaves the barrel. What is the average impulse imparted to the bullet?


(a)    1.8 N-s


(b)    Zero


        (c)    9 N-s


        (d)    0.9 N-s


Grade:upto college level

1 Answers

Simran Bhatia
10 years ago

(d)
Given F = 600-(2 × 105t)
The force is zero at time t, given by.
0 = 600 - 2 × 105 t
? t =  600/(2×?10?^5 )=3×?10?^(-3)  second
? Impulse = ∫_0^t¦? Fdt  = ∫_0^(3×?10?^(-3))¦? (600 -2×?10?^5 t)  dt??
= [600 t -(2×?10?^5 t^2)/2]^(3×?10?^(-3) )
= 600×3×?10?^(-3)- ?10?^5 (3×?10?^(-3) )^2
=1.8-0.9 =0.9 Ns